# IPAD: Inverse Prompt for AI Detection – A Reliable and Explainable LLM-Generated Essay Detector

### v2

@@ 3.3 Interpretability Assessment of IPAD
To assess the explainability improvement of IPAD, we designed an interpretability assessment with ten participants evaluating one HWT and one LGT article. We used IPAD version 2 due to its superior OOD performance and attack resistance. Participants compared three online detection platforms [^6] https://www.scribbr.com/ai-detector/ [^7] https://quillbot.com/ai-content-detector [^8] https://app.gptzero.me/ with IPAD’s process (which displayed input texts, predicted prompts, regenerated texts, and final judgments). After evaluation, participants rated IPAD on four key explainability dimensions. Transparency received strong ratings (40%:5, 60%:4), with participants appreciating the visibility of intermediate processes. Trust scores were more varied (10%:3, 70%:4, 20%:5), but IPAD was generally considered more convincing than single-score detectors. Satisfaction was mixed (30%:3, 30%:4, 40%:5), with participants acknowledging better detection but raising concerns about energy efficiency since IPAD runs three LLMs. Debugging received unanimous 5s, as participants could easily analyze the predicted prompt and regenerated text to verify the decision-making process. If needed, users could refine the generated content by adjusting instructions, such as specifying a word count, making IPAD a more effective and user-friendly tool compared to black-box detectors. We further analyzed the different linguistic features of HWT prompts and LGT prompts as illustrated in Appendix  G.
@@ 4.1 AI detectors Methods and challenges
AI text detection methods can be broadly categorized into four approaches Wu et al. (2025): watermarking, statistics-based methods, neural-based methods, and human-assisted methods.
Watermarking technology inserts specific patterns into training datasets Shevlane et al. (2023); Gu et al. (2022) or manipulates the model output during inference to embed a watermark Lucas and Havens (2023). However, watermarking needs to access the LLM deployment and can face attacks, such as identifying and erasing the watermark Hou et al. (2024).  Statistics-based methods analyze inherent textual features to identify language patterns Kalinichenko et al. (2003); Hamed and Wu (2023), but their effectiveness depends on corpus size and model diversity Wu et al. (2025). Some other statistical methods use n-gram probability divergence Yang et al. (2024b) or similarity between original and revised texts Mao et al. (2024); Zhu et al. (2023) while still face robustness challenges under adversarial attacks Wu et al. (2025)Neural-based methods such as RoBERTa Liu et al. (2020), Bert Devlin et al. (2019), and XLNet Yang et al. (2019) have been robust in domain-specific tasks. Adversarial learning techniques are increasingly being used Yang et al. (2024a) to increase effectiveness in attacked datasets.
In addition to automated methods, human involvement plays a key role in detecting AI-generated text Wu et al. (2025).  Human-assisted detection leverages human intuition and expertise to identify inconsistencies such as semantic errors and logical flaws that may not be easily caught by algorithms Uchendu et al. (2023); Dugan et al. (2023). Moreover, given the challenges of current AI detection tools, which often lack verifiable evidence Chaka (2023), human involvement becomes even more critical to ensure the reliable and explainable detection.
@@ 4.2 Prompt Inverter techniques and applications
Prompt extraction techniques aim to reverse-engineer the prompts that generate specific outputs from LLMs. Approaches include black-box methods like output2prompt Zhang et al. (2024a), which extracts prompts based on model outputs without access to internal data, and logit-based methods like logit2prompt Mitka (2024), which rely on next-token probabilities but are constrained by access to logits. Adversarial methods can bypass some defenses but are model-specific and fragile Zhang et al. (2024c). Despite the success of some zero-shot LLM-inversion based methods Li and Klabjan (2024); Yu et al. (2024), they are mostly naive usage of prompting LLMs, which makes them poor in prompt extraction accuracy and robustness.
@@ 5 Conclusion
This paper introduces  IPAD (Inverse Prompt for AI Detection), a framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and two  Distinguishers that examines how well the input texts align with the predicted prompts. One is the  Prompt-Text Consistency Verifier (PTCV) which evaluates direct alignment between predicted prompts and input text, and the other is  Regeneration Comparator (RC) that examines content similarity by comparing input texts with the corresponding regenerated texts. Empirical evaluations demonstrate that IPAD outperforms the strongest baselines by 9.05% (Average Recall) on in-distribution data, 12.93% (AUROC) on out-of-distribution (OOD) data, and 5.48% (AUROC) on attacked data. The combination of the two modules suggests that combining self-consistency checks of generative models with multi-step reasoning for evidential explainability holds promise for future AI detection systems in real-world scenarios. An interpretability assessment reveals that IPAD enhances trust and transparency by allowing users to examine decision-making evidence.
Limitations
@@ Appendix A Fine-tune Dataset
  • Instructions-2M Morris et al. (2024), a collection of 2 million user prompts and system prompts, from which we used 30,000 prompts.
  • ShareGPT Zhang et al. (2024d), an open platform where users share ChatGPT prompts and responses, from which we used 500 samples.
  • Unnatural Instructions Zhang et al. (2024d), a dataset of creative instructions generated by OpenAI’s models, from which we used 500 samples.
  • OUTFOX dataset Koike et al. (2024), which contains 15,400 essay problem statements, student-written essays, and LLM-generated essays.
The first three datasets aims to enhance the general querying capability of the Prompt Inverter, and are all released under the MIT license. All the samples we used are the same to the samples randomly selected in  Zhang et al. (2024a). The last dataset aims to enhance the familiarity of the Prompt Inverter with the data of the essay to detect the LLM-generated essays, and are created and examined by  Koike et al. (2024), We specifically used the LLM-generated essays and problem statements for this supervised fine-tuning (SFT). There are 45,400 training pairs in total.
Given that essay data are diverse, we utilize only the OUTFOX dataset Koike et al. (2024). To adapt this dataset for training our Distinguisher, we enhance it to align with the model’s requirements. The original dataset consists of 14,400 training triplets of essay problem statements, student-written essays, and LLM-generated essays. To further process the data, we apply the Prompt Inverter to both student-written and LLM-generated essays, generating corresponding Predicted Prompts. These Predicted Prompts are then used to regenerate texts via ChatGPT, i.e. gpt-3.5-turbo. Following this procedure, we construct a total of 28,800 training samples, with an equal distribution of positive and negative examples (14,400 each).
@@
Instruction "What is the prompt $P$ that generates the Input Text $T$?" "Can LLM generate the input text $T$ through the prompt $P$?" "$T^{\prime}$ is generated by LLM, determine whether $T$ is also generated by LLM with a similar prompt."
@@
Appendix B Complete Algorithm
The complete inference pipeline is summarized in Algorithm 1.
Algorithm 1 IPAD Detection Procedure
2:Prediction $\hat{Y}\in\{\mathrm{HWT},\mathrm{LGT}\}$ and evidence $\mathcal{E}$
3:$P\leftarrow\mathbf{f}_{\text{inv}}(T)$triangleright Inverse-prompt prediction
4:$T^{\prime}\leftarrow\mathbf{f}_{\text{LLM}}(P)$triangleright Regenerate text using P
5:$\boldsymbol{z}^{\text{PTCV}}\leftarrow\mathbf{f}_{\text{PTCV}}(P,T)$
6:$p_{\text{PTCV}}\leftarrow\dfrac{\sigma(\boldsymbol{z}^{\text{PTCV}}_{\text{yes}})}{\sigma(\boldsymbol{z}^{\text{PTCV}}_{\text{yes}})+\sigma(\boldsymbol{z}^{\text{PTCV}}_{\text{no}})}$
7:$\boldsymbol{z}^{\text{RC}}\leftarrow\mathbf{f}_{\text{RC}}(T^{\prime},T)$
8:$p_{\text{RC}}\leftarrow\dfrac{\sigma(\boldsymbol{z}^{\text{RC}}_{\text{yes}})}{\sigma(\boldsymbol{z}^{\text{RC}}_{\text{yes}})+\sigma(\boldsymbol{z}^{\text{RC}}_{\text{no}})}$
9:$\hat{p}\leftarrow w\cdot p_{\text{PTCV}}+(1-w)\cdot p_{\text{RC}}$
10:if $\hat{p}>\tau$ then
11:     $\hat{Y}\leftarrow\mathrm{LGT}$
12:else
13:     $\hat{Y}\leftarrow\mathrm{HWT}$
14:end if
15:$\mathcal{E}\leftarrow\bigl{(}P,\;p_{\text{PTCV}},\;p_{\text{RC}},\;\hat{p}\bigr{)}$
16:return $(\hat{Y},\;\mathcal{E})$
Appendix C AUROC formula
Since our model predicts binary labels, we follow the  Wilcoxon-Mann-Whitney statistic Calders and Jaroszewicz (2007) to calculate the Area Under Receiver Operating Characteristtic curve (AUROC):
@@ Appendix C AUROC formula
Appendix D Ablation study data structures
@@
Appendix E DPIC (decouple prompt and intrinsic characteristics) Prompt Extraction Zero-shot Prompts Yu et al. (2024)
@@
Appendix F Comparison with DPIC
@@
Appendix G Different Linguistic Features of HWT prompts and LGT prompts
This subsection of the evaluation aims to explore the linguistic features of prompts generated by HWT and LGT through the Prompt Inverter. We analyzed 1000 samples generated by HWT and 1000 samples generated by LGT, which are randomy selected from both in-distribution data and OOD.
The analysis is first conducted using the Linguistic Feature Toolkik (lftk) [^9] https://lftk.readthedocs.io/en/latest/, a commonly used general-purpose tool for linguistic features extraction, which provides a total of 220 features for text analysis. Upon applying this toolkit, we identified 20 features with significant differences in average values between the two groups, out of which 3 features showed statistically significant differences with p-values less than 0.05. These 3 differences can be summarized as one main aspects:  syntactic complexity. Beyond these, we referred to the LIWC framework  [^10] https://www.liwc.app/, which defines 7 function words variables and 4 summary variables. By comparing the difference, two of these 11 features is significantly distinguishable:  the pronoun usage and  the level of analytical thinking.
One of the primary distinctions between the HWT prompts and the LGT prompts is sentence complexity. LGT prompts are typically more complex, characterized by longer sentence lengths (mean value of 1.514 and 1.794), higher syllable counts (mean values of total syllabus three are 1.572 and 3.042), and more stop-words (mean values of 9.88 and 10.045). HWT prompts, on the other hand, are characterized by shorter, less complex sentences that are easier to process and understand.
Beyond the differences in syntactic complexity, we also explored variables in LIWC. We did the difference comparison by using HWT and LGT prompts as inputs for ChatGPT, for example, instructing with the prompts ’determine the pronoun usage of this sentence, answer first person, second person, or third person’ and ’determine the level of analytical thinking of these sentences, answer a number from 1 to 5’. The results show that there are distinguish difference in pronoun usage and analytical thinking level. The HWT prompts frequently use second-person pronouns (e.g., ’you’) - 75 occurrences per 1,000 prompts - due to the subjective tone often employed in HWT. In contrast, LGT prompts primarily feature first- and third-person pronouns, with second-person pronouns appearing only 2 per 1,000 prompts. LGT prompts typically present instructions and questions in a more objective manner. LGT prompts show higher analytical thinking levels than HWT prompts. With level 1 as the lowest and level 5 as the highest, LGT has 68.9% of level 4 and 24.3% of level 5, but HWT has only 48.0% of level 4, and 0.8% of level 5. It suggests that LGT prompts encourage more analytical thinking, while HWT prompts tend to focus more on concrete examples, with less emphasis on critical analysis.
Appendix H IPAD and DPIC prompt inverter examples

### v3

@@ 3.3 Interpretability Assessment of IPAD
To assess the explainability improvement of IPAD, we designed an interpretability assessment with ten participants evaluating one HWT and one LGT article. We used IPAD version 2 due to its superior OOD performance and attack resistance. Participants compared three online detection platforms (i.e., Scribbr, QuillBot, GPTZero) with IPAD’s process (which displayed input texts, predicted prompts, regenerated texts, and final judgments). After evaluation, participants rated IPAD on four key explainability dimensions. Transparency received strong ratings (40%:5, 60%:4), with participants appreciating the visibility of intermediate processes. Trust scores were more varied (10%:3, 70%:4, 20%:5), but IPAD was generally considered more convincing than single-score detectors. Satisfaction was mixed (30%:3, 30%:4, 40%:5), with participants acknowledging better detection but raising concerns about energy efficiency since IPAD runs three LLMs. Debugging received unanimous 5s, as participants could easily analyze the predicted prompt and regenerated text to verify the decision-making process. If needed, users could refine the generated content by adjusting instructions, such as specifying a word count, making IPAD a more effective and user-friendly tool compared to black-box detectors.
@@ 4.1 AI detectors Methods and challenges
Recent studies have explored diverse strategies for detecting AI-generated text. Watermarking embeds identifiable patterns during training r28; r27 or inference r29, but requires model access and is vulnerable to erasure attacks r30. Statistics-based methods treat output distributions as detection signals. DetectGPT r35 and Fast-DetectGPT bao2023fast locate LGT in regions of negative curvature of log-probability; Lastde xu2025trainingfree and Glimpse bao2025glimpse exploit token-probability dynamics and partial-distribution prediction. Other statistical approaches rely on n-gram divergence or revision similarity r32; r31; r38; r37; r36, though robustness remains limited r12. Regeneration-based methods compare model rewrites with originals: RAIDAR r37 and MAGRET huang-etal-2025-magret observe stronger edits on human text; DNA-GPT r36 and TOCSIN r11 measure continuation or deletion-based differences. Neural approaches fine-tune large encoders (e.g., RoBERTa r42, BERT r44, XLNet r45) with adversarial or contrastive objectives r14; r46, while human-in-the-loop methods provides complementary semantic judgment and explainability r22; r50; r49.
@@ 4.2 Prompt Inverter techniques and applications
Prompt extraction techniques aim to reverse-engineer the prompts that generate specific outputs from LLMs. Approaches include black-box methods like output2prompt r4, which extracts prompts based on model outputs without access to internal data, and logit-based methods like logit2prompt r51, which rely on next-token probabilities but are constrained by access to logits. Adversarial methods can bypass some defenses but are model-specific and fragile r52. Despite the success of some zero-shot LLM-inversion based methods r53; r62, they are mostly naive usage of prompting LLMs, which makes them poor in prompt extraction accuracy and robustness.
@@ 5 Conclusion
This paper introduces  IPAD (Inverse Prompt for AI Detection), a framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and two  Distinguishers that examines how well the input texts align with the predicted prompts. One is the  Prompt-Text Consistency Verifier (PTCV) which evaluates direct alignment between predicted prompts and input text, and the other is  Regeneration Comparator (RC) that examines content similarity by comparing input texts with the corresponding regenerated texts. Empirical evaluations demonstrate that IPAD outperforms the strongest baselines by 9.05% (Average Recall) on in-distribution data, 12.93% (AUROC) on out-of-distribution (OOD) data, and 5.48% (AUROC) on attacked data. IPAD also performs robustly on structured datasets. While the local alignment in RC approach provides explicit interpretability, it is more sensitive to adversarial attacks. In contrast, the global distribution in PTCV matching approach implicitly learns generative LLM’s distributional properties, which offers more robustness while maintaining explainability. The combination of the two modules suggests that combining self-consistency checks of generative models with multi-step reasoning for evidential explainability holds promise for future AI detection systems in real-world scenarios. An interpretability assessment reveals that IPAD enhances trust and transparency by allowing users to examine decision-making evidence. Overall, IPAD establishes a new paradigm for more robust, reliable, and interpretable AI detection systems to combat the misuse of LLMs.
@@
Acknowledgments
This work was supported in part by NSFC grants (62432008 and 623B2002), in part by RGC RIF grant R6021-20, in part by RGC TRS grant T43-513/23N-2, in part by RGC CRF grants (C7004-22G, C1029-22G and C6015-23G), in part by NSFC/RGC grant CRS_HKUST601/24, and in part by RGC GRF grants (16207922, 16207423 and 16203824).
@@ Appendix A Fine-tune Dataset
  • Instructions-2M r5, a collection of 2 million user prompts and system prompts, from which we used 30,000 prompts.
  • ShareGPT r6, an open platform where users share ChatGPT prompts and responses, from which we used 500 samples.
  • Unnatural Instructions r6, a dataset of creative instructions generated by OpenAI’s models, from which we used 500 samples.
  • OUTFOX dataset r3, which contains 15,400 essay problem statements, student-written essays, and LLM-generated essays.
The first three datasets aims to enhance the general querying capability of the Prompt Inverter, and are all released under the MIT license. All the samples we used are the same to the samples randomly selected in  r4. The last dataset aims to enhance the familiarity of the Prompt Inverter with the data of the essay to detect the LLM-generated essays, and are created and examined by  r3, We specifically used the LLM-generated essays and problem statements for this supervised fine-tuning (SFT). There are 45,400 training pairs in total.
Given that essay data are diverse, we utilize only the OUTFOX dataset r3. To adapt this dataset for training our Distinguisher, we enhance it to align with the model’s requirements. The original dataset consists of 14,400 training triplets of essay problem statements, student-written essays, and LLM-generated essays. To further process the data, we apply the Prompt Inverter to both student-written and LLM-generated essays, generating corresponding Predicted Prompts. These Predicted Prompts are then used to regenerate texts via ChatGPT, i.e. gpt-3.5-turbo. Following this procedure, we construct a total of 28,800 training samples, with an equal distribution of positive and negative examples (14,400 each).
@@
Instruction "What is the prompt $P$ that generates the Input Text TT?" "Can LLM generate the input text $T$ through the prompt PP?" $T^{\prime}$ is generated by LLM, determine whether $T$ is also generated by LLM with a similar prompt."
@@
Appendix B AUROC formula
Since our model predicts binary labels, we follow the  Wilcoxon-Mann-Whitney statistic r56 to calculate the Area Under Receiver Operating Characteristtic curve (AUROC):
@@ Appendix B AUROC formula
Appendix C Ablation study data structures
@@
Appendix D DPIC (decouple prompt and intrinsic characteristics) Prompt Extraction Zero-shot Prompts r62
@@
Appendix E Comparison with DPIC
@@
Appendix F IPAD and DPIC prompt inverter examples
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We explained the contributions listed in the abstract and introduction in Section 2-4.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We present the limitation in the section Conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: There is no theorem or lemma in this paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The codes are presented in the anonymous github, and the results can be reproduced by following the readme.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The codes are presented in the anonymous github, and the results can be reproduced by following the readme. The dataset for this paper are publicly available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental details are presented.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We set the temperature parameter of the large language models to 0, so the results are deterministic.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We use one A800-80GB gpu and one Nvidia v100 to conduct our experiments.
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code of ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: This paper has no ethics problem.
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [Yes]
  48. Justification: We discuss the potential societal impacts mainly in section 3.2.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: There are no such risks in our paper.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: : We have cited the original papers and included proper license.
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: We do not release new assets.
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and research with human subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: We have no crowdsourcing experiments and human subjects.
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: We have no crowdsourcing experiments and human subjects.
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  75. 16.
    Declaration of LLM usage
  76. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  77. Answer: [Yes]
  78. Justification: Pretrained LLMs are used as backbones in our method, which is clearly stated in this paper.
  79. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.